[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jan 6, 2016

Mark DeMaria - NOAA Federal mark.demaria at noaa.gov
Wed Jan 6 13:12:12 UTC 2016


Here is today's update on NHC's computer operations.

Mark

-----
TAFB:

Martin had problems with gfe2nawips script Tue evening. The problem was
caused by a full disk on the compute server, which Matt fixed. The disk was
filled up by a process being used to create training material. That script
is being checked so this problem does not  reoccur.

HSU:

None.


Other Issues:

The blue rack containing the old sound system will be removed this week.
TSB will coordinate with NCO on a replacement sound system.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160106/a02ea72b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list